OpenStudy (anonymous):

3^x=9^x+1 how to solve

OpenStudy (turingtest):

\[3^x=9^{x+1}\]?

OpenStudy (anonymous):

Is it??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you show formulas or/and your work please

OpenStudy (anonymous):

You can write it as: \[3^x = (3^2)^{x+1}\] \[\large 3^x = 3^{2x+2}\] Now exponents must be equal.. So equate the exponents and find x from there..

OpenStudy (anonymous):

so do you cancel the threes and solve for x when you have ^x=2x+2

OpenStudy (anonymous):

See, the two equations are equal.. Their base(3) is also the same.. If they are equal then their exponents must be equal.. So we here equate the exponents..

OpenStudy (anonymous):

In general: \[\huge x^a = x^b \implies \color{green}{a = b}\]

OpenStudy (anonymous):

I'm still confused I'm not just looking for the answer but I'm better at analyzing the work what values would a and b be

OpenStudy (anonymous):

Here a = x and b = 2x + 2

OpenStudy (anonymous):

so x=2x+2

OpenStudy (anonymous):

so i subtract from from both sides then subtract x so i get x=-2

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

See first his reply.. He replied yes to Turning test..

OpenStudy (anonymous):

oh... so this problem is easier than what i have then....

OpenStudy (anonymous):

Ha ha ha.. Yes it is..